HanAkoSolo (jamierox4ev3r):

Check my work please. Not sure if I'm going in the right direction here

HanAkoSolo (jamierox4ev3r):

\[\lim_{h \rightarrow 0}\left( \frac{ \sqrt{9+h}-3 }{ h } \right)\]

HanAkoSolo (jamierox4ev3r):

I assume that i would have to multiply both numerator and denominator with a conjugate

OpenStudy (anonymous):

correct

HanAkoSolo (jamierox4ev3r):

and that conjugate would be \(\large\sqrt{9+h}+3\)

HanAkoSolo (jamierox4ev3r):

alright. At least that assumption was correct

HanAkoSolo (jamierox4ev3r):

so from there, since i would multiply \(\sqrt{9+h}-3\times \sqrt{9+h}+3\)

HanAkoSolo (jamierox4ev3r):

since \(\sqrt{x^{2}}\) is equal to \(x\), then \(\sqrt{9+h}\times \sqrt{9+h}\) would be equal to \(9+h\)

HanAkoSolo (jamierox4ev3r):

Since the middle terms cancel out, I believe we are left with \(\Large\lim_{h \rightarrow 0} \frac{h}{h(\sqrt{9+h} +3}\)

OpenStudy (anonymous):

in the numerator 9+h-9 in the denominator ? you are correct.

HanAkoSolo (jamierox4ev3r):

Nice! So I wasn't sure if I was multiplying things correctly oh, and I kind of forgot to close the parentheses in the denominator, but other that, it looks good to you? @surjithayer

HanAkoSolo (jamierox4ev3r):

\(\Large\lim_{h \rightarrow 0} \frac{\color{red}h}{\color{red}h(\sqrt{9+h} +3}\) The things in red will cancel out, so you should be left with \(\Large\lim_{h \rightarrow 0} \frac{1}{\sqrt{9+h} +3}\)

HanAkoSolo (jamierox4ev3r):

@nincompoop from here, would I just plug in zero in place of h? (assuming, of course, that I hadn't taken any erroneous steps)

OpenStudy (aravindg):

I didnt read the entire steps above but if the denominator doesnt go to 0 you can plug in 0 for h.

HanAkoSolo (jamierox4ev3r):

at this point, it doesn't.

HanAkoSolo (jamierox4ev3r):

took a lot of simplifying, which I pray I got correct XD. Thanks for the tip, movie buddy :3

OpenStudy (nincompoop):

assuming that you've done correctly up to the last portion, all you have to do is plug in the value of h = 0 and simplify

OpenStudy (nincompoop):

you need to show me the original function tho

HanAkoSolo (jamierox4ev3r):

alright. I've already verified that most of my steps were correct with surji, and here is the original fxn \(\Large\lim_{h \rightarrow 0}\left( \frac{ \sqrt{9+h}-3 }{ h } \right)\)

HanAkoSolo (jamierox4ev3r):

@AravindG I think I already outlined that somewhere up there :P

OpenStudy (jhannybean):

You got this :) You just need to stay confident!

HanAkoSolo (jamierox4ev3r):

thanks :) @Jhannybean

OpenStudy (aravindg):

Yeah you are right. The limit exists.

OpenStudy (jhannybean):

\[\boxed{\lim_{h \rightarrow 0}\left( \frac{ \sqrt{9+h}-3 }{ h } \right) = \frac{1}{6}}\]

OpenStudy (nincompoop):

there are few key points that you probably already learned whether or not limit may or may not exist based on the function itself without doing all of the work. it cuts many of your work

OpenStudy (nincompoop):

thanks, jhan

OpenStudy (nincompoop):

I knew there was something odd LAUGHING OUT LOUD

OpenStudy (nincompoop):

jhan :*

HanAkoSolo (jamierox4ev3r):

Alright. Good to know. \(\Large\lim_{h \rightarrow 0} \frac{1}{\sqrt{9+h} +3}\) \(\Large\frac{1}{\sqrt{9+0} +3} = \frac{1}{3+3}\)

OpenStudy (jhannybean):

Yep.

HanAkoSolo (jamierox4ev3r):

and that would give you \(\Large\frac{1}{6}\). Alright then, that makes sense!

HanAkoSolo (jamierox4ev3r):

So at that point, you can plug in 0, since it wouldn't become undefined. g'nice

OpenStudy (nincompoop):

ye why is that?

OpenStudy (nincompoop):

why can't we just plug it in immediately?

HanAkoSolo (jamierox4ev3r):

because then the whole fraction would be set over 0. As long as the denominator is zero, we conclude that the limit does not exist, when in reality it does

OpenStudy (nincompoop):

interesting

HanAkoSolo (jamierox4ev3r):

At least I think that's what the prof. said o-o lol

OpenStudy (nincompoop):

so what does it mean that the limit = 1/6 ?

HanAkoSolo (jamierox4ev3r):

why dost thou question me profusely? XD

OpenStudy (nincompoop):

fundamental concepts and analysis

HanAkoSolo (jamierox4ev3r):

I see o-o

OpenStudy (jhannybean):

That means at some value of x, the function will approach y at 1/6,

HanAkoSolo (jamierox4ev3r):

^that's the technical meaning of a limit

OpenStudy (nincompoop):

and then what?

HanAkoSolo (jamierox4ev3r):

At this point, I'm not sure what you're asking

OpenStudy (nincompoop):

haha

HanAkoSolo (jamierox4ev3r):

Seriously XD I can't keep track of all these questions. What more do you want than the technical definition? o-o

OpenStudy (nincompoop):

just telling me what the formal definition of limit does not tell me what you understand about it, that's all

HanAkoSolo (jamierox4ev3r):

Something to do with reimann sum right? :P lol

HanAkoSolo (jamierox4ev3r):

Anyhow, thank you, I'll close this for now :)

HanAkoSolo (jamierox4ev3r):

@nincompoop Is the reimann sum the same as the epsilon-delta limit definition?! Because if it is, then we barely went over that in class. And if they are indeed the same thing, then I do understand what you're saying

OpenStudy (nincompoop):

they are related

HanAkoSolo (jamierox4ev3r):

I see